OpenStudy (anonymous):

Solve by elimination. 5x + 6y = 1 and 10x + 12 = 2. What is the ordered pair?

OpenStudy (anonymous):

-1,1

OpenStudy (saifoo.khan):

No Solution.

OpenStudy (anonymous):

no solution??

OpenStudy (saifoo.khan):

these equations have no solutions.

OpenStudy (dumbcow):

is the 2nd equation 10x+12y + 2 ????

jimthompson5910 (jim_thompson5910):

There are an infinite number of solutions to the system of equations 5x + 6y = 1 and 10x + 12y = 2

OpenStudy (anonymous):

oh my bad!!! the second equation IS 10x + 12y = 2

OpenStudy (anonymous):

oh, no solution then

OpenStudy (anonymous):

I aplogize

OpenStudy (saifoo.khan):

Still, No Solution.

OpenStudy (saifoo.khan):

np.

OpenStudy (anonymous):

So the answer of an infinite number of solutions is not correct? It is NO solutions?

jimthompson5910 (jim_thompson5910):

There are an infinite number of solutions because the second equation is really the first equation but each term is multiplied by 2. So effectively, the two equations form the exact same line.

OpenStudy (anonymous):

There are infinite many solutions.

OpenStudy (dumbcow):

jim_thompson is correct, there are infinite solutions multiply 1st equation by -2, then add the equations -10x -12y = -2 10x +12y = 2 -------------- 0 =0

OpenStudy (saifoo.khan):

not infinity.

OpenStudy (saifoo.khan):

these are parallel lines, if we graph them

jimthompson5910 (jim_thompson5910):

Check your graph again, you should only get one line (basically one line is directly on top of the other)

OpenStudy (saifoo.khan):

its NOT.

jimthompson5910 (jim_thompson5910):

show me the graph if you can

jimthompson5910 (jim_thompson5910):

along with what you're inputting

OpenStudy (anonymous):

Let me phrase it again correctly so there is no confusion. 5x + 6y = 1 and 10x + 12y = 2

OpenStudy (anonymous):

There are infinite many solutions.

jimthompson5910 (jim_thompson5910):

Yep, those equations are basically the same one. So there are an infinite number of solutions. For instance, say x=-1 and y=1. Plug these into the first equation to get 5(-1)+6(1)=1 -5+6=1 1=1 Now plug them into the second equation 10(-1)+12(1)=2 -10+12=2 2=2 Since that solution clearly works, this means that there is at least one solution. This one solution is (-1,1). You can keep going forever showing there are more solutions.

OpenStudy (anonymous):

x = 6 n+5, y = -5 n-4, This is what x and y equal. You can put any number in for n and it will be correct